# Large Language Model Compression with Global Rank and Sparsity Optimization

### v1

@@ 6 Experiments
We compare our method with several state-of-the-art techniques for sparsification and composite approximation, all of which aim to compress pre-trained language models: SparseGPT Frantar and Alistarh [2023] is a second-order pruning method for LLMs that solves a layer-wise reconstruction problem. WANDA Sun et al. [2023] prunes weights based on their estimated importance using activation statistics. BESA Xu et al. [2024] is a recent pruning method that allocates sparsity across layers in a differentiable manner to minimize overall pruning error. LPAF Ren and Zhu [2023] first applies first-order unstructured pruning to obtain a low-rank sparse model. Then, sparsity-aware SVD is used to decompose the sparse matrices into a low-rank form $\mathbf{A}\mathbf{B}$. Finally, mixed-rank fine-tuning is used to retrain $\mathbf{A}\mathbf{B}$. LoSparse Li et al. [2023] first initializes a low-rank matrix $\mathbf{A}\mathbf{B}$ and a residual sparse matrix $\mathbf{S}$ via SVD. Then, all parameters are iteratively updated while applying structured pruning to $\mathbf{S}$.
6.1 Main Results
In Table 1, we present the results of layer-wise Bernoulli distribution-based probabilistic pruning applied after the global RPCA decomposition of each model. Our method consistently outperforms previous unstructured pruning approaches across both LLaMA and LLaMA-2 models at a 50% compression ratio (the reduction in total parameters from both the low-rank and sparse components relative to the original model), demonstrating the effectiveness of our composite approximation strategy. Notably, the memory requirements of the LoSparse method exceed those of full-parameter fine-tuning, rendering it impractical to test on LLaMA models; comparisons with LoSparse Li et al. [2023] using the DeBERTaV3-base He et al. [2023] model can be found in Appendix B.
From a time and memory efficiency perspective, our approach primarily incurs computational costs during the decomposition and probabilistic pruning stages. The RPCA decomposition is performed directly on each matrix independently, requiring no data and incurring minimal memory and time costs (e.g., decomposing all matrices in the 7B model over ten iterations requires only a few minutes). During the probabilistic pruning optimization, only a small sample of data is required to compute forward activations for updating the probability distributions, making our memory consumption comparable to methods like WANDA, and significantly more efficient than LoSparse.
Global pruning consistently outperforms layer-wise pruning, which in turn is superior to single-matrix pruning. In terms of perplexity, global pruning achieves an average reduction of 0.2 compared to layer-wise pruning across all models. We recommend applying probabilistic pruning following global RPCA decomposition for more robust results, as this approach preserves structural consistency during policy gradient estimation.
Perplexity Changes During Compressing
Starting from the initial layer, we sequentially decomposed and pruned each layer of the LLaMA2-7B model, with the resulting performance changes shown in Figure 3. The performance degradation generally correlates with the number of pruned layers; however, an unexpected performance boost occurs when the last layer is compressed, underscoring the importance of maintaining structural consistency within the model. Interestingly, decomposing and pruning the first layer led to a slight improvement over the original model’s performance, suggesting that early layers may indeed contain some redundant information.
Figure 4: The rank distribution of low-rank matrices for each module after pruning.
Low-Rank Component Rank Distribution
We analyzed the rank distribution of low-rank components in each module after probabilistic pruning. Figure 4 shows these ranks along with the layer-wise averages. Modules with higher ranks in the low-rank component tend to have sparser counterparts in the sparse component to meet the compression target. The general trend of increasing rank in deeper layers suggests that redundancy varies across the network, with later layers capturing more complex representations, underscoring the need for careful pruning in these layers. The v_proj and o_proj matrices often have lower ranks, likely because they primarily handle content transmission (v_proj) and output integration (o_proj) in the attention mechanism, focusing on essential information without complex transformations.
6.2 Ablation Studies
Table 2: Effect of RPCA iterations on model performance. Here, “Avg. Rank" denotes the average rank of the low-rank component, and “Sparsity" represents the sparsity of the sparse component.
Iter Sparsity Avg. Rank PPL $\downarrow$ Avg. Error
1 0.8080 936.56 32497 0.6248
3 0.4123 2109 5.18 0.0243
10 0.4005 2217 5.13 0.0213
100 0.4090 2198 5.16 0.0209
Effect of RPCA Iterations
We investigated the effect of RPCA iterations on the performance of the LLaMA2-7B model. Table 2 shows the performance results for different numbers of RPCA iterations, with no additional pruning applied post-decomposition. The results indicate that only a few RPCA iterations are needed to achieve an effective decomposition, providing a solid starting point for subsequent pruning. Additionally, the “Avg. Error" column represents the average approximation error for each matrix, offering insight into the model’s tolerance to error. Similar to findings in the quantization field, large models exhibit robustness to approximation errors. Finally, the fact that performance even surpasses the original model after decomposition suggests the presence of some redundant parameters in the pre-trained model.
Heuristic Threshold-Based Pruning
We also conducted experiments using a heuristic threshold-based approach after RPCA decomposition. In this method, we prune the low-rank component $\mathbf{L}$ by setting singular values below a specific threshold to zero and remove low-magnitude elements from the sparse component $\mathbf{S}$ without applying probabilistic masking or additional optimization. We explored various thresholds and sparsity levels to observe their effects on the model’s average rank, sparsity, and perplexity. Table 3 summarizes the results for LLaMA2-7B under different threshold and sparsity configurations. The results indicate that while the low-rank component is slightly more critical than the sparse component, the sparse component still cannot be entirely removed. When only the low-rank component is retained, the model experiences a performance collapse; similarly, retaining only the sparse component leads to a similar collapse. This explains why performance degradation is often severe when low-rank approximation or sparse approximation is used in isolation.
Table 3: Effect of heuristic threshold-based pruning on model performance. Singular values below the threshold are set to zero.
Threshold Avg. Rank Sparsity PPL $\downarrow$
0.5 1342 0.6 5.84
1 684 0.6 11.14
2 214 0.6 2909.51
Max 0 0.6 NaN
0.5 1342 0.8 7.63
0.5 1342 1.0 NaN
6.3 Discussion
Our method, leveraging composite approximation with policy gradient optimization, effectively reduces model size while maintaining high performance. Although our experiments focus on LLaMA models, the approach is model-agnostic and applicable to various architectures as it operates entirely at the matrix level. Additional results at higher compression ratios are provided in Appendix D, where our method continues to outperform baseline pruning techniques despite more pronounced performance degradation. This demonstrates that, while robust, the method faces practical limits to compression without significant performance loss. Beyond a certain threshold, model collapse becomes unavoidable. Although fine-tuning can alleviate such degradation, our work emphasizes compression efficacy without any parameter updates. For hyperparameter selection, such as the penalty parameter $\lambda$ in RPCA, we recommend using default settings. In our experiments, dynamic updates to these hyperparameters yielded no substantial benefits, likely due to RPCA’s inherent convergence properties. To further enhance the efficiency of composite approximation, we plan to explore methods for integrating low-rank and sparse components, such as reshaping matrices via structured pruning. Our approach holds potential for broader application.
7 Conclusion
This work aims to solve the compression problem of pre-trained large language models and proposes a two-stage low-rank and sparse composite approximation compression method. First, the weight matrix is decomposed into low-rank subspace and sparse subspace through RPCA, which significantly reduces the search space; then, the global probability distribution optimization technology based on Bernoulli sampling is used to automatically identify and retain the most important low-rank and sparse components. Compared with the traditional method of manually setting thresholds, this scheme can adaptively allocate the rank and sparsity of different layers, and can achieve better reasoning performance and robustness on various benchmarks without large-scale backpropagation or fine-tuning. In the future, further research can be conducted on the combination with technologies such as quantization and knowledge distillation and their application in larger-scale models to provide a more efficient solution for the deployment of large models in multiple scenarios.
lmpact Statement
This paper presents work whose goal is toadvance the field of Machine Learning. There aremany potential societal consequences of ourwork, none which we feel must be specificalyhighlighted here.
Appendix A Knowledge Neurons
Transformer-based architectures, particularly large language models (LLMs), serve as repositories of linguistic and factual knowledge Geva et al. [2021], Dai et al. [2022]. This knowledge is intricately distributed across the network’s feed-forward networks (FFNs) and attention mechanisms, forming the basis for accurate language understanding and generation. Figure 5 provides an illustrative depiction of how such knowledge is encoded, stored, and attributed across these components, with factual information such as "Ireland’s capital is Dublin" encapsulated through complex interactions.
A.1 Impact of Pruning FFN Layers
FFNs act as key-value storage within Transformer models, encoding linguistic and factual information as neuron activations Geva et al. [2021]. Specific neurons, often referred to as knowledge neurons, are responsible for capturing and representing precise knowledge. For example, one neuron may activate to encode “Q27: Ireland," while its interplay with others encodes the factual relationship “Capital: Dublin."
Pruning FFN layers introduces the following risks:
A.2 Effect of Pruning Attention Mechanisms
  • Knowledge Attribution: Attention mechanisms identify and link related entities, such as establishing the factual connection between “Ireland" and “Dublin."
Pruning attention mechanisms poses distinct challenges:
  • Impaired Attribution: Removing attention heads or weights can disrupt critical connections between tokens, such as the association between “Ireland" and “Dublin," resulting in factual inconsistencies.
A.3 Challenges in Simultaneous Pruning
A.4 Proposed Mitigation Strategies
A.5 Conclusion
Appendix B Comparison with LoSparse
LoSparse is an iterative pruning method that achieves the target compression ratio by progressively pruning and fine-tuning both the low-rank and sparse components. This iterative procedure ensures strong performance by continuously adjusting the pruned model’s parameters during compression. In contrast, our proposed method, CAP (Composite Approximation Pruning), allows for flexible pruning configurations with significantly reduced fine-tuning requirements. Specifically, fine-tuning is only applied to the compressed model, thereby reducing the computational and resource overhead compared to LoSparse.
Method MNLI (%) QNLI (%) RTE (%)
w/o pruning 90.5 94.0 82.0
LoSparse 81.55 81.42 53.43
CAP1 70.07 78.13 52.71
CAP2 72.04 78.13 53.79
CAP3 82.21 82.54 53.43
CAP4 82.75 85.52 57.04
Table 4: Zero-shot accuracy results on MNLI, QNLI, and RTE for the pruned LLaMA-7B model at 80% parameter retention. CAP1 uses RPCA with heuristic thresholds, CAP2 combines RPCA and policy gradient optimization, CAP3 uses RPCA with full parameter fine-tuning, and CAP4 combines RPCA, policy gradient optimization, and full parameter fine-tuning. The results show the best performance for CAP4.
Hyper-parameter Value
Batch size 64
Learning rate 3e-5
Table 5: Hyper-parameter setup for pruning DeBERTaV3-base using LoSparse with float32 precision.
Analysis: As shown in Table 4, our method achieves competitive or superior performance compared to LoSparse under the same compression ratio. CAP4 demonstrates the highest accuracy across all tasks, surpassing LoSparse while requiring significantly fewer resources for fine-tuning.
Specifically:
  • CAP1 and CAP2 Performance: CAP1 (RPCA with heuristic thresholds) and CAP2 (RPCA + policy gradient optimization) achieve reasonable accuracy but fall slightly short of LoSparse due to the lack of parameter fine-tuning. This is expected, as LoSparse iteratively prunes and fine-tunes the low-rank and sparse components during compression.
  • CAP3 and CAP4 Performance: CAP3 (RPCA + full parameter fine-tuning) improves performance, matching or slightly exceeding LoSparse on certain tasks. CAP4, which combines RPCA, policy gradient optimization, and fine-tuning, consistently outperforms LoSparse across all tasks, achieving the best results.
Efficiency: Unlike LoSparse, which fine-tunes parameters iteratively during pruning, CAP fine-tunes only after compression. This strategy significantly reduces computational costs and memory usage. Since fine-tuning is performed on a compressed model, the resource consumption is much lower than that of LoSparse, making CAP an efficient alternative for resource-constrained scenarios.
Conclusion: The results highlight the effectiveness of our method in maintaining high accuracy while reducing the computational overhead associated with iterative pruning and fine-tuning. By leveraging RPCA for decomposition and policy gradient optimization for adaptive pruning, CAP achieves state-of-the-art performance in compressed LLMs with minimal resource requirements.
Appendix C Reproduction Report
This appendix presents a comprehensive report detailing the reproduction process for our RPCA-based model compression methodology, incorporating policy gradient optimization. The implementation builds upon WANDA’s code base and integrates components from additional open-source libraries, to which we extend our gratitude.
C.1 Implementation Overview
C.2 Code Structure
The implementation is organized into three main components:
C.3 Running the Pruning Process
To reproduce our results, follow these steps:
  1. 1.
    Environment Setup:
    • Ensure Python 3.8 or later is installed.
    • Install the necessary dependencies:
              pip install torch \
                transformers \
                numpy \
                tqdm \
                matplotlib \
                json \
                argparse
  2. 2.
    Execution:
      bash main.sh
C.4 Key Implementation Details
C.5 Results
C.6 Conclusion
Methods BoolQ RTE HellaSwag Winogrande ARC-e ARC-c OBQA
w/o Pruning 75.08 66.09 56.94 69.93 75.25 41.89 34.60
Magnitude 42.23 52.35 25.86 48.38 26.64 21.50 14.00
SparseGPT 60.86 52.71 29.10 51.78 33.08 17.58 13.40
WANDA 37.83 51.99 26.78 49.41 28.79 19.54 13.20
CAP (75%) 63.33 55.71 31.67 54.01 35.73 22.88 16.12
Methods BoolQ RTE HellaSwag Winogrande ARC-e ARC-c OBQA
w/o Pruning 77.71 62.82 57.14 68.90 76.39 43.52 31.40
Magnitude 43.49 49.10 25.69 50.83 26.09 20.73 16.00
SparseGPT 60.34 54.15 30.28 53.12 33.75 20.39 14.20
WANDA 38.22 52.17 26.95 50.51 28.03 19.62 13.20
CAP (75%) 62.67 56.15 32.11 55.39 36.12 23.02 16.37
Appendix D Comparison under Higher Sparsity
We evaluate our proposed method (CAP) against three other pruning methods—Magnitude, SparseGPT, and WANDA—on two versions of the LLaMA-7B model under a parameter retention rate of 75%. The results are presented in Tables 6 and 7.
Observations:
  • Magnitude and SparseGPT: These methods show noticeable performance degradation under 75% parameter retention, especially on tasks that require factual reasoning (e.g., OBQA) or commonsense understanding (e.g., HellaSwag).
  • WANDA: WANDA performs slightly better than SparseGPT on certain datasets but is generally less competitive compared to CAP, highlighting the advantages of CAP’s probabilistic pruning mechanism.
  • CAP excels in challenging settings: By leveraging RPCA for decomposition and policy gradient optimization for adaptive pruning, CAP is able to selectively retain the most informative parameters, ensuring superior performance even under extreme sparsity.
Conclusion: The results demonstrate that CAP offers a robust and efficient approach to model compression. By eliminating heuristic thresholds and adopting fine-grained pruning strategies, CAP surpasses existing methods while maintaining computational simplicity and avoiding the need for post-pruning fine-tuning.
Appendix E Comparison with LPAF
In this section, we compare our proposed CAP method with LPAF in terms of performance on GLUE tasks using the BERT-base model, considering three different compression ratios: 50%, 25%, and 16% of the original parameters. LPAF (Low-rank Prune-and-Factorize) applies structured pruning followed by sparsity-aware SVD decomposition and mixed-rank fine-tuning. In contrast, CAP leverages Robust Principal Component Analysis (RPCA) for adaptive decomposition and probabilistic pruning with policy gradient optimization. The main advantage of CAP lies in its automatic selection of compression configurations without requiring extensive fine-tuning.

### v3

@@ 3 Experiments
We compare our approach with several compression techniques: SparseGPT (Frantar and Alistarh, 2023) is a second-order pruning method for LLMs that solves a layer-wise reconstruction problem. WANDA (Sun et al., 2024) prunes weights based on their estimated importance using activation statistics. OATS (Zhang and Papyan, 2024) performs optimal sparsity allocation across transformer layers using second-order information. OWL (Yin et al., 2023) and AlphaPruning (Lu et al., 2024) are layer-wise allocation methods that optimize sparsity distribution. SLiM (Mozaffari et al., 2024) combines low-rank approximation with sparsity and quantization, featuring probabilistic quantization error fitting. LPAF (Ren and Zhu, 2023) first applies first-order unstructured pruning to obtain a low-rank sparse model. Then, sparsity-aware SVD is used to decompose the sparse matrices into a low-rank form $\mathbf{A}\mathbf{B}$. Finally, mixed-rank fine-tuning is used to retrain $\mathbf{A}\mathbf{B}$. We also compare against dense SVD-based methods SVD-LLM2 (Wang et al., 2025c), Dobi-SVD (Wang et al., 2025b), Basis Sharing (Wang et al., 2025a) and structured pruning method LoSparse (Li et al., 2023).
3.1 Comparison with Unstructured Pruning Methods
We compare CAP with recent unstructured pruning methods across multiple large language models. Table 1 presents a comprehensive comparison including both uniform sparsity methods (SparseGPT, Wanda, DSNoT, OATS) and layerwise allocation methods (OWL, AlphaPruning) at 30%, 40%, and 50% compression ratios. Note that OWL and AlphaPruning are layerwise allocation methods that optimize sparsity distribution across layers, and we implement them using Wanda as the base pruning method for fair comparison. CAP consistently achieves competitive or superior performance across different model architectures and sizes.
3.2 Extension to Modern Instruction-Tuned Models
To validate CAP’s effectiveness on the latest generation of models, we evaluated LLaMA-3.1-8B-Instruct and Qwen2.5-7B at 50% sparsity. We report on Chain-of-Thought reasoning (GSM8K) and long-context understanding (LongBench-v2), where standard pruning methods struggle.
Table 2: Performance on Modern LLMs (50% Sparsity). Left: LLaMA-3.1-8B-Instruct on reasoning and long-context tasks. Right: Qwen2.5-7B on diverse benchmarks.
LLaMA-3.1-8B-Instruct Qwen2.5-7B
Method GSM8K LongBench-v2 WikiText Task SparseGPT Wanda CAP Dense
(8-shot, %) (Avg, %) (PPL) (50%) (50%) (50%) (0%)
Dense 84.5 30.4 6.01 NarrativeQA 4.95 16.30 18.52 31.9
Wanda (50%) 45.6 25.1 7.26 GovReport 33.52 32.13 34.05 53.4
CAP (50%) 56.8 27.2 6.61 Lcc (Code) 39.70 45.14 46.92 58.1
Improvement +11.2 +2.1 -0.65 TriviaQA 89.35 88.70 89.85 92.5
As shown in Table 2, CAP significantly outperforms Wanda on challenging tasks. For LLaMA-3.1-Instruct, CAP recovers +11.2% accuracy on GSM8K, suggesting that the preserved low-rank backbone is crucial for maintaining the precise reasoning circuits disrupted by unstructured pruning.
3.3 Comparison with Joint Compression Methods
Since methods like LoSparse are based on structured pruning and require extensive retraining, we compare CAP with SLiM, a state-of-the-art method that jointly applies quantization, sparsity, and low-rank approximation. We also include comparisons with other joint compression approaches including JSQ (Guo et al., 2024), a joint sparsity and quantization method that optimizes sparsity and quantization parameters simultaneously, and L2QER (Zhang et al., 2024a), which combines low-rank decomposition, quantization, and sparsity in a sequential manner.
While both SLiM and CAP structurally combine low-rank and sparse components, their technical approaches differ fundamentally: SLiM primarily focuses on using low-rank decomposition to fit quantization errors through probabilistic reformulation and numerical integration to find optimal quantization parameters, whereas CAP focuses on the synergy between low-rank and sparse decomposition through RPCA, where the low-rank component emerges from joint optimization rather than serving as an error fitting tool. Table 3 presents the comparison on representative models at 50% unstructured sparsity. The results demonstrate that CAP consistently outperforms existing joint compression methods across different model sizes and architectures.
Table 3: Comparison at 50% unstructured sparsity. Zero-shot accuracy (%) on representative models. LoRA variants: Naive-LoRA uses basic error compensation; SLiM-LoRA incorporates weight salience; SLiM-LoRAQ additionally quantizes the adapter.
Method Quantization OPT LLaMA-2
1.3B 2.7B 6.7B 13B 7B 13B
Dense - 43.4 45.5 48.3 48.7 56.6 60.8
Magnitude Group AbsMax 32.1 39.9 36.4 32.3 47.0 51.0
SparseGPT OPTQ 38.7 43.4 47.0 47.4 51.1 55.9
Wanda OPTQ 41.0 42.9 46.5 46.8 53.6 56.8
JSQ JSQ 38.9 35.5 42.8 30.7 52.3 57.0
L2QER Group AbsMax 38.4 41.3 45.1 OOM 50.6 OOM
Naive-LoRA QuantizationW 40.4 43.4 46.6 47.3 51.5 55.3
SLiM-LoRA QuantizationW 41.9 43.5 47.1 48.0 54.3 57.9
SLiM-LoRAQ QuantizationW 41.7 43.6 47.2 47.9 54.2 57.3
CAP (Ours) OPTQ 41.7 44.8 48.2 48.3 55.1 59.2
3.4 Comprehensive Comparison on GLUE Tasks
We evaluate CAP on downstream tasks using the GLUE benchmark with BERT-base. Table 4 compares CAP against various compression paradigms including pre-training distillation, task-specific distillation, structured pruning, and matrix factorization methods.